# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ 4.1.3. Generative Learning in CF
  • Generative Learning Paradigm in self-supervised learning CF has evolved from variational autoencoding to masked autoencoding and denoised diffusion. (i) Variational Autoencoding. Different VAE-based methods in self-supervised learning CF adopt various perspectives to improve the model. For instance, CVAE (Li and She, 2017) and MD-CVAE (Zhu and Chen, 2022) incorporate item content information to regulate the collaborative latent variable modeling. RecVAE (Shenbin et al., 2020) follows the general paradigm in Mult-VAE (Liang et al., 2018) but introduces a novel composite prior distribution and a new approach to setting hyperparameters. MacridVAE (Ma et al., 2019) and CaD-VAE (Wang et al., 2023a) focus on modeling the latent factor behind interactions to capture the latent variable distribution. BiVAE (Truong et al., 2021) extends the VAE paradigm to symmetrically generate user-item interactions. FastVAE (Chen et al., 2022a) enhances model efficiency with inverted multi-index. SE-VAE (Cho and Oh, 2022) extends latent variable modeling with multi-experts and stochastic expert selection. (ii) Masked Autoencoding. Both AutoCF (Xia et al., 2023) and GFormer (Li et al., 2023e) adopt the mask-reconstruction paradigm, where they automatically identify valuable edges (user-item interactions) in the graph and reconstruct them using the learned user/item representations. (iii) Denoised Diffusion. DiffRec (Wang et al., 2023d) has introduced the diffusion model into collaborative filtering. This method leverages diffusion to denoise perturbed user ratings or latent variables, which helps mitigate the high resource costs associated with large-scale item prediction.
  • Generation Target. Both VAE-based methods (Liang et al., 2018; Ma et al., 2019; Shenbin et al., 2020; Chen et al., 2022a; Cho and Oh, 2022; Wang et al., 2023a) and diffusion-based methods (Wang et al., 2023d) share a similar auto-encoding architecture. In these methods, the user ratings to items, typically represented as vectors, serve as the generation target for the model. However, some VAE-based methods like CVAE and MD-CAE also generate item content to achieve hybrid generation. BiVAE adopts a symmetric approach by modeling the user and item latent variables, with user-item interactions as the generation target. On the other hand, in graph-based methods (Xia et al., 2023; Li et al., 2023e), valuable edges in the graph are masked and then reconstructed.
@@ 4.1.4. Adversarial Learning in CF
  • Adversarial Learning Paradigm in collaborative filtering encompasses both non-differentiable and differentiable training approaches. IRGAN (Wang et al., 2017) introduced the idea of GANs into recommendation systems, sampling item indices from the generator for discrimination. However, the discrete nature of these item indices prevents the gradient from the discriminator from propagating to the generator for optimization. To overcome this issue, IRGAN employs policy gradient techniques, which are also adopted by Krishnan et al. (Krishnan et al., 2018). In the context of differentiable adversarial training, several methods such as CFGAN (Chae et al., 2018), RAGAN (Chae et al., 2019), and Conv-GCF (Yuan et al., 2020b) feed the generated continuous real-valued variables into the discriminator. On the other hand, ABinCF (Wang et al., 2019b) and AugCF (Wang et al., 2019c) use the Gumbel trick to approximate sampling, enabling end-to-end adversarial training.
@@ 4.2. Sequential Recommendation
CoSeRec (Liu et al., 2021a) 2021 ArXiv Data-based Natural InfoNCE-based
@@
ContrastVAE (Wang et al., 2022e) 2022 CIKM Model-based Natural InfoNCE-based
EC4SRec (Wang et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
MCLSR (Wang et al., 2022b) 2022 CIKM Model-based Natural InfoNCE-based
@@
TCPSRec (Tian et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
@@
GCL4SR (Zhang et al., 2022a) 2022 IJCAI Data & Model-based Natural InfoNCE-based
@@
ICLRec (Chen et al., 2022b) 2022 WWW Data & Model-based Natural InfoNCE-based
@@
FDSA_CL (Hao et al., 2023a) 2023 TKDE Model-based Natural InfoNCE-based
ContraRec (Wang et al., 2023b) 2023 TOIS Data-based Natural InfoNCE-based
@@
IOCRec (Li et al., 2023c) 2023 WSDM Data & Model-based Natural InfoNCE-based
DCRec (Yang et al., 2023b) 2023 WWW Model-based Natural InfoNCE-based
Meta-SGCL (Hao et al., 2023b) 2024 ICDE Model-based Natural InfoNCE-based
@@
GL Method Venue Year Generative Learning Paradigm Generation Target
@@
PTUM (Wu et al., 2020b) 2020 EMNLP Masked Autoencoding Behavior in Sequence
PeterRec (Yuan et al., 2020a) 2020 SIGIR Masked Autoencoding Item in Sequence
@@
VSAN (Zhao et al., 2021b) 2021 ICDE Variational Autoencoding Item in Sequence
@@
ContrastVAE (Wang et al., 2022e) 2022 CIKM Variational Autoencoding Item in Sequence
DiffuASR (Liu et al., 2023b) 2023 CIKM Denoised Diffusion Augmented Item Sequence
Diff4Rec (Wu et al., 2023b) 2023 MM Denoised Diffusion User-Item Interaction
@@
AOS4Rec (Zhao et al., 2021a) 2020 IJCAI Non-differentiable Item
@@
SAO (Manotumruksa and Yilmaz, 2020) 2020 SIGIR Differentiable Feature of Sequence
@@ 4.2.2. Contrastive Learning in SeqRec
  • View Creation in SeqRec includes data-based, feature-based, and model-based methods. (i) Data-based View Creation. Given that the input data is in a sequential format, numerous methods employ sequence augmentation techniques to generate diverse views of the input item sequence. In general, early works (Zhou et al., 2020; Xie et al., 2022b; Liu et al., 2021a; Shin et al., 2021; Li et al., 2021) propose non-heuristic and random augmentation methods including Crop, Mask, Reorder, Shuffle, Substitute, and Insert to obtain different input sequences as different views. TiCoSeRec (Dang et al., 2023) goes a step further by incorporating temporal information and introducing five operators (e.g., Ti-Crop and Ti-Reorder). These operators are designed to transform the input sequence into a uniform representation while taking into account the variations in time intervals. In EC4SRec (Wang et al., 2022a), importance scores are calculated for each item in a sequence, which is then utilized to guide heuristic and explanation-driven operations for generating informative views. In addition to these methods, several data-based approaches consider constructing additional data resources such as item-item transition graphs (Zhang et al., 2022a; Yang et al., 2023b), item co-interaction (Yang et al., 2023b), and user sequences for each item (Lin et al., 2022a) to obtain diverse viewpoints. (ii) Feature-based View Creation. Feature-based methods employ various augmentation on the encoded features for view creation. For instance, ACVAE (Xie et al., 2021) utilizes variational auto-encoders to encode latent features of item sequences. It then shuffles these latent features to generate different negative views of the sequence. Additionally, MIC (Nie et al., 2022) applies random dropout on the encoded features, resulting in two distinct views of user/item representation for contrastive learning. (iii) Model-based View Creation. Model-based view creation in SeqRec encompasses two distinct approaches. The first approach involves constructing specific neural modules to encode views of interest, such as intent-aware representation (Ma et al., 2020; Chen et al., 2022b; Qin et al., 2024), hypergraph representation (Xia et al., 2021b), item attributes representation (Hao et al., 2023a), variational Transformer (Hao et al., 2023b), long-short term representations (Liu et al., 2024b) and utilizing graph neural networks and to encode embeddings from auxiliary graphs (Wang et al., 2022b; Yang et al., 2023b; Lin et al., 2022a). The second approach involves applying techniques such as random dropout (Cheng et al., 2021; Qiu et al., 2021a; Wang et al., 2022e; Qiu et al., 2022) on the model parameters to generate different outputs as distinct views.
@@
  • Pair Sampling in SeqRec has both natural sampling and score-based heuristic sampling. (i) Natural Sampling. In most cases in SeqRec, natural sampling methods (Zhou et al., 2020; Liu et al., 2021a; Bian et al., 2021; Xie et al., 2022b; Du et al., 2022a; Wang et al., 2022a; Zhang et al., 2022a) naturally considering the generated views from the same item sequences as positive views and otherwise negative views. In addition, there are other natural pairing relations utilized in SeqRec. For instance, methods like SSI (Yuan et al., 2021), TCPSRec (Tian et al., 2022a), and DCRec (Yang et al., 2023b) construct positive pairs using various combinations, such as pairing a single item or a sub-sequence with the whole sequence it belongs to. Besides, ICSRec (Qin et al., 2024) uses clustering to generate intent prototypes and aligns the intent view with related prototypes for fine-grained intent CL. (ii) Score-based Sampling. Compared to natural sampling, score-based sampling utilizes necessary calculate to determine the positive pair. COTREC (Xia et al., 2021a) combines the last-clicked item with the predicted next item to form a positive pair. H2SeqRec (Li et al., 2021), DuoRec (Qiu et al., 2022), and MrTransformer (Ma et al., 2023) employ sequence overlap as a metric to determine positive pairs. MIC (Nie et al., 2022) first applies k-means clustering to group samples, and then samples within the same group form positive pairs. Similarly, MISS (Guo et al., 2022) employs item embedding distance as a metric to sample positive pairs with varying probabilities.
@@
4.2.3. Generative Learning in SeqRec
@@
  • Generative Learning Paradigm in SeqRec predominantly follows the approach of mask-autoencoding. For instance, BERT4Rec (Sun et al., 2019) initially attempts to randomly mask an item within the item sequence. Then, it proceeds with the encoding process using a transformer and aims to reconstruct the masked item. Given that using transformers for modeling item sequence data (Kang and McAuley, 2018) in SeqRec is a prevalent choice, the natural approach of mask-autoencoding with transformers has become a primary form in subsequent works (Wu et al., 2020b; Yuan et al., 2020a; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Liu et al., 2021b; Du et al., 2022a). Furthermore, MAERec (Ye et al., 2023b) adopts the concept of mask-autoencoding from graph data (Hou et al., 2022b; Li et al., 2022c) to enhance the representation learning of items. Another notable approach in generative modeling is seen in SVAE (Sachdeva et al., 2019), which leverages variational auto-encoding as a backbone. This method utilizes recurrent neural network to encode sequential records into latent variables that follow the normal distribution and subsequently generate the next-item with a re-parametrization trick (Kingma and Welling, 2013b). This line of work is further extended by methods like VSAN (Zhao et al., 2021b) and ContrastVAE (Wang et al., 2022e) by incorporating self-attention modules or contrastive learning on latent variables. Recently, denoised diffusion has emerged as a promising paradigm in sequential recommendation (SeqRec) due to its strong generation capabilities. Both DiffuASR (Liu et al., 2023b) and Diff4Rec (Wu et al., 2023b) leverage diffusion generation to create high-quality data, which in turn enhances the training of recommendation models through data augmentation.
@@
  • Generation Target in SeqRec with generative learning exhibits a diverse range of approaches. In general, several works (Sachdeva et al., 2019; Yuan et al., 2020a; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Zhao et al., 2021b; Liu et al., 2021b; Du et al., 2022a; Wang et al., 2022e) follow the initial step of BERT4Rec, which involves masking and reconstructing the items within the sequence. PTUM considers the behavior sequence of users as the target for generation, aiming to effectively model user behavior patterns for recommendation purposes. On the other hand, U-BERT focuses on generating user reviews for items to capture user behaviors and enhance recommendations, particularly in content-insufficient domains. Recently, MAERec has incorporated the concept of mask-autoencoding from the graph domain. It utilizes this approach to generate paths in the item transition graph, effectively pre-training item embeddings for improved sequential recommendation performance. For diffusion-based methods, DiffuASR generates augmented item sequences to enrich sparse data for model training, while Diff4Rec pre-trains a diffusion model by corrupting and reconstructing user-item interactions in the latent space, which are then used to produce diverse augmentations for sparse user-item interactions.

### v3

@@ 4.1.3. Reconstructive Learning in CF
  • Reconstructive Learning Paradigm in self-supervised learning CF has evolved from variational autoencoding to masked autoencoding and denoised diffusion. (i) Variational Autoencoding. Different VAE-based methods in self-supervised learning CF adopt various perspectives to improve the model. For instance, CVAE (Li et al., 2017) and MD-CVAE (Zhu et al., 2022) incorporate item content information to regulate the collaborative latent variable modeling. RecVAE (Shenbin et al., 2020) follows the general paradigm in Mult-VAE (Liang et al., 2018) but introduces a novel composite prior distribution and a new approach to setting hyperparameters. MacridVAE (Ma et al., 2019) and CaD-VAE (Wang et al., 2023c) focus on modeling the latent factor behind interactions to capture the latent variable distribution. BiVAE (Truong et al., 2021) extends the VAE paradigm to symmetrically generate user-item interactions. FastVAE (Chen et al., 2022a) enhances model efficiency with inverted multi-index. SE-VAE (Cho et al., 2022) extends latent variable modeling with multi-experts and stochastic expert selection. (ii) Masked Autoencoding. Both AutoCF (Xia et al., 2023) and GFormer (Li et al., 2023b) adopt the mask-reconstruction paradigm, where they automatically identify valuable edges (user-item interactions) in the graph and reconstruct them using the learned user/item representations. (iii) Denoised Diffusion. DiffRec (Wang et al., 2023e) has introduced the diffusion model into collaborative filtering. This method leverages diffusion to denoise perturbed user ratings or latent variables, which helps mitigate the high resource costs associated with large-scale item prediction.
  • Reconstruction Target. Both VAE-based methods (Liang et al., 2018; Ma et al., 2019; Shenbin et al., 2020; Chen et al., 2022a; Cho et al., 2022; Wang et al., 2023c) and diffusion-based methods (Wang et al., 2023e) share a similar auto-encoding architecture. In these methods, the user ratings to items, typically represented as vectors, serve as the reconstruction target for the model. However, some VAE-based methods like CVAE and MD-CAE also generate item content to achieve hybrid generation. BiVAE adopts a symmetric approach by modeling the user and item latent variables, with user-item interactions as the reconstruction target. On the other hand, in graph-based methods (Xia et al., 2023; Li et al., 2023b), valuable edges in the graph are masked and then reconstructed.
@@ 4.1.4. Adversarial Learning in CF
  • Adversarial Learning Paradigm in collaborative filtering encompasses both non-differentiable and differentiable training approaches. IRGAN (Wang et al., 2017) introduced the idea of GANs into recommendation systems, sampling item indices from the generator for discrimination. However, the discrete nature of these item indices prevents the gradient from the discriminator from propagating to the generator for optimization. To overcome this issue, IRGAN employs policy gradient techniques, which are also adopted by Krishnan et al. (Krishnan et al., 2018). In the context of differentiable adversarial training, several methods such as CFGAN (Chae et al., 2018), RAGAN (Chae et al., 2019), and Conv-GCF (Yuan et al., 2020a) feed the generated continuous real-valued variables into the discriminator. On the other hand, ABinCF (Wang et al., 2019a) and AugCF (Wang et al., 2019b) use the Gumbel trick to approximate sampling, enabling end-to-end adversarial training.
@@
4.1.5. In-depth Analysis of SSL-based Collaborative Filtering
The three main SSL techniques-contrastive learning, reconstructive learning, and adversarial learning-offer distinct advantages and challenges in CF. Contrastive learning enhances model robustness by leveraging diverse views from user-item interactions (Wu et al., 2021a; Zhou et al., 2023c), but its effectiveness can be hindered by the quality of constructed views. Reconstructive learning, through methods like VAEs and MAEs, excels in reconstructing interactions for richer representations (Shenbin et al., 2020; Xia et al., 2023), though it may face scalability issues in large datasets. Adversarial learning improves recommendation accuracy by using a discriminator to optimize outputs (Wang et al., 2017; Chae et al., 2018), yet its non-differentiable nature can complicate training for CF. Ultimately, the choice of technique should align with the specific needs of the collaborative filtering task, considering factors like data availability and computational resources.
@@ 4.2. Sequential Recommendation
CoSeRec (Liu et al., 2021d) 2021 ArXiv Data-based Natural InfoNCE-based
@@
ContrastVAE (Wang et al., 2022d) 2022 CIKM Model-based Natural InfoNCE-based
EC4SRec (Wang et al., 2022b) 2022 CIKM Data-based Natural InfoNCE-based
MCLSR (Wang et al., 2022e) 2022 CIKM Model-based Natural InfoNCE-based
@@
TCPSRec (Tian et al., 2022b) 2022 CIKM Data-based Natural InfoNCE-based
@@
GCL4SR (Zhang et al., 2022c) 2022 IJCAI Data & Model-based Natural InfoNCE-based
@@
ICLRec (Chen et al., 2022d) 2022 WWW Data & Model-based Natural InfoNCE-based
@@
FDSA_CL (Hao et al., 2023) 2023 TKDE Model-based Natural InfoNCE-based
ContraRec (Wang et al., 2023a) 2023 TOIS Data-based Natural InfoNCE-based
@@
IOCRec (Li et al., 2023g) 2023 WSDM Data & Model-based Natural InfoNCE-based
DCRec (Yang et al., 2023a) 2023 WWW Model-based Natural InfoNCE-based
Meta-SGCL (Hao et al., 2024) 2024 ICDE Model-based Natural InfoNCE-based
@@
LS4SRec (He et al., 2025) 2025 TOIS Model-based Natural Explicit
@@
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
@@
PTUM (Wu et al., 2020a) 2020 EMNLP Masked Autoencoding Behavior in Sequence
PeterRec (Yuan et al., 2020b) 2020 SIGIR Masked Autoencoding Item in Sequence
@@
VSAN (Zhao et al., 2021a) 2021 ICDE Variational Autoencoding Item in Sequence
@@
ContrastVAE (Wang et al., 2022d) 2022 CIKM Variational Autoencoding Item in Sequence
DiffuASR (Liu et al., 2023a) 2023 CIKM Denoised Diffusion Augmented Item Sequence
Diff4Rec (Wu et al., 2023c) 2023 $M$ Denoised Diffusion User-Item Interaction
@@
AOS4Rec (Zhao et al., 2021b) 2020 IJCAI Non-differentiable Item
@@
SAO (Manotumruksa et al., 2020) 2020 SIGIR Differentiable Feature of Sequence
@@ 4.2.2. Contrastive Learning in SeqRec
  • View Creation in SeqRec includes data-based, feature-based, and model-based methods. (i) Data-based View Creation. Given that the input data is in a sequential format, numerous methods employ sequence augmentation techniques to generate diverse views of the input item sequence. In general, early works (Zhou et al., 2020; Xie et al., 2022b; Liu et al., 2021d; Shin et al., 2021; Li et al., 2021) propose non-heuristic and random augmentation methods including Crop, Mask, Reorder, Shuffle, Substitute, and Insert to obtain different input sequences as different views. TiCoSeRec (Dang et al., 2023) goes a step further by incorporating temporal information and introducing five operators (e.g., Ti-Crop and Ti-Reorder). These operators are designed to transform the input sequence into a uniform representation while taking into account the variations in time intervals. In EC4SRec (Wang et al., 2022b), importance scores are calculated for each item in a sequence, which is then utilized to guide heuristic and explanation-driven operations for generating informative views. In addition to these methods, several data-based approaches consider constructing additional data resources such as item-item transition graphs (Zhang et al., 2022c; Yang et al., 2023a), item co-interaction (Yang et al., 2023a), and user sequences for each item (Lin et al., 2022a) to obtain diverse viewpoints. (ii) Feature-based View Creation. Feature-based methods employ various augmentation on the encoded features for view creation. For instance, ACVAE (Xie et al., 2021) utilizes variational auto-encoders to encode latent features of item sequences. It then shuffles these latent features to generate different negative views of the sequence. Additionally, MIC (Nie et al., 2022) applies random dropout on the encoded features, resulting in two distinct views of user/item representation for contrastive learning. (iii) Model-based View Creation. Model-based view creation in SeqRec encompasses two distinct approaches. The first approach involves constructing specific neural modules to encode views of interest, such as intent-aware representation (Ma et al., 2020; Chen et al., 2022d; Qin et al., 2024), hypergraph representation (Xia et al., 2021b), item attributes representation (Hao et al., 2023), variational Transformer (Hao et al., 2024), long-short term representations (Liu et al., 2024b) and utilizing graph neural networks and to encode embeddings from auxiliary graphs (Wang et al., 2022e; Yang et al., 2023a; Lin et al., 2022a). The second approach involves applying techniques such as random dropout (Cheng et al., 2021; Qiu et al., 2021a; Wang et al., 2022d; Qiu et al., 2022) on the model parameters to generate different outputs as distinct views.
@@
  • Pair Sampling in SeqRec has both natural sampling and score-based heuristic sampling. (i) Natural Sampling. In most cases in SeqRec, natural sampling methods (Zhou et al., 2020; Liu et al., 2021d; Bian et al., 2021; Xie et al., 2022b; Du et al., 2022a; Wang et al., 2022b; Zhang et al., 2022c) naturally considering the generated views from the same item sequences as positive views and otherwise negative views. In addition, there are other natural pairing relations utilized in SeqRec. For instance, methods like SSI (Yuan et al., 2021), TCPSRec (Tian et al., 2022b), and DCRec (Yang et al., 2023a) construct positive pairs using various combinations, such as pairing a single item or a sub-sequence with the whole sequence it belongs to. Besides, ICSRec (Qin et al., 2024) uses clustering to generate intent prototypes and aligns the intent view with related prototypes for fine-grained intent CL. (ii) Score-based Sampling. Compared to natural sampling, score-based sampling utilizes necessary calculate to determine the positive pair. COTREC (Xia et al., 2021a) combines the last-clicked item with the predicted next item to form a positive pair. H2SeqRec (Li et al., 2021), DuoRec (Qiu et al., 2022), and MrTransformer (Ma et al., 2023) employ sequence overlap as a metric to determine positive pairs. MIC (Nie et al., 2022) first applies k-means clustering to group samples, and then samples within the same group form positive pairs. Similarly, MISS (Guo et al., 2022) employs item embedding distance as a metric to sample positive pairs with varying probabilities.
@@
4.2.3. Reconstructive Learning in SeqRec
@@
  • Reconstructive Learning Paradigm in SeqRec predominantly follows the approach of mask-autoencoding. For instance, BERT4Rec (Sun et al., 2019) initially attempts to randomly mask an item within the item sequence. Then, it proceeds with the encoding process using a transformer and aims to reconstruct the masked item. Given that using transformers for modeling item sequence data (Kang and McAuley, 2018) in SeqRec is a prevalent choice, the natural approach of mask-autoencoding with transformers has become a primary form in subsequent works (Wu et al., 2020a; Yuan et al., 2020b; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Liu et al., 2021b; Du et al., 2022a). Furthermore, MAERec (Ye et al., 2023b) adopts the concept of mask-autoencoding from graph data (Hou et al., 2022; Li et al., 2023c) to enhance the representation learning of items. Another notable approach in generative modeling is seen in SVAE (Sachdeva et al., 2019), which leverages variational auto-encoding as a backbone. This method utilizes recurrent neural network to encode sequential records into latent variables that follow the normal distribution and subsequently generate the next-item with a re-parametrization trick (Kingma and Welling, 2013). This line of work is further extended by methods like VSAN (Zhao et al., 2021a) and ContrastVAE (Wang et al., 2022d) by incorporating self-attention modules or contrastive learning on latent variables. Recently, denoised diffusion has emerged as a promising paradigm in sequential recommendation (SeqRec) due to its strong generation capabilities. Both DiffuASR (Liu et al., 2023a) and Diff4Rec (Wu et al., 2023c) leverage diffusion generation to create high-quality data, which in turn enhances the training of recommendation models through data augmentation.
@@
  • Reconstruction Target in SeqRec with generative learning exhibits a diverse range of approaches. In general, several works (Sachdeva et al., 2019; Yuan et al., 2020b; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Zhao et al., 2021a; Liu et al., 2021b; Du et al., 2022a; Wang et al., 2022d) follow the initial step of BERT4Rec, which involves masking and reconstructing the items within the sequence. PTUM considers the behavior sequence of users as the target for generation, aiming to effectively model user behavior patterns for recommendation purposes. On the other hand, U-BERT focuses on generating user reviews for items to capture user behaviors and enhance recommendations, particularly in content-insufficient domains. Recently, MAERec has incorporated the concept of mask-autoencoding from the graph domain. It utilizes this approach to generate paths in the item transition graph, effectively pre-training item embeddings for improved sequential recommendation performance. For diffusion-based methods, DiffuASR generates augmented item sequences to enrich sparse data for model training, while Diff4Rec pre-trains a diffusion model by corrupting and reconstructing user-item interactions in the latent space, which are then used to produce diverse augmentations for sparse user-item interactions.